Using the expanded accounting equation, calculate and enter the answers for each question. You will need to use the answers you calculate for beginning and ending equity to answer the rest of the questions.

Assets Liabilities
Beginning of Year: $29,000 $16,000
End of Year: $64,000 $29,000

1) What is the equity at the beginning of the year?

Beginning Equity:13000

2) What is the equity at the end of the year?

Ending Equity: 35000

3) If the owner contributes $8,200 and the owner withdraws $39,600, how much is net income (loss)?

Net Income (Loss):

4) If net income is $3,200 and owner withdrawals are $7,400, how much did the owner contribute (owner, capital)?

Owner, Capital:

5) If the owner contributes $16,000 and net income is $19,600, how much did the owner withdraw (owner, withdrawals)?

Owner, Withdrawals:

6) If the owner contributes $44,500 and the owner withdraws $3,900, how much is net income (loss)?

Net Income (Loss):
Answer #1

1) Beginning equity = 29000-16000 = 13000

2) Ending equity = 64000-29000 = $35000

3) Net income (loss) = 13000+8200-39600-35000 = $53400

4) Owner's capital = 13000+3200-7400-35000 = 26200

5) Owner's withdrawal = 13000+16000+19600-35000 = 13600

6) Net income (loss) = 13000+44500-3900-35000 = -18600